Q1 the accuracy in the measurement of the diameter of hydrogen atom as 1.06*10^-10m is

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
accuracy = delta d/d = 0.01 * 10^-10/ 1.06 * 10^-10 = 1/106
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
Khimraj
4 years ago
accuracy = \Delta d/d = 0.01 * 10^-10/ 1.06 * 10^-10 = 1/106
….........................................................................................................